Web Analytics

Julio is playing a card game. The probability that an ace card is drawn is 0.08. The probability that a black card is drawn is 0.5. The probability that a black ace card is drawn is 0.04.

Which equation shows the correct use of the addition rule to determine the probability that Julio will draw an ace card or a black card?

A0.08−0.04=0.04
B0.5+0.04−0.08=0.46
C0.08+0.5−0.04=0.54
D0.08+0.5=0.58
E0.5+0.04=0.54